A "Quiescent IOP" (IOP 6) from noon today through noon tomorrow; Persistent fog radiosounde plan. 

We spent the morning adding thermocouples. In order of time:

  • SP - 945a
  • SH - 10a
  • MH - 10:39a  

Swapped out the thermocouple that was already there. This was a different type, so we swapped it anyway to be the same type as all the others. Put old one the box #2.

  • PC - 11:20a
  • UP - 11:39a
  • LC - 12:07p 
    • John pushed the wires into the sonic more tightly - thinking there was a loose wire connection. Will monitor to see if that fixed it.

CC - 1230p

DC - 12:47p

MW - 1:04p

ISS: Launched the ~2p and ~415p. 

  • No labels